\begin{thebibliography}{56}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Ashurst et~al.(2022)Ashurst, Hine, Sedille, and
  Carlier}]{ashurst2022ai}
Carolyn Ashurst, Emmie Hine, Paul Sedille, and Alexis Carlier. 2022.
\newblock Ai ethics statements: Analysis and lessons learnt from neurips
  broader impact statements.
\newblock In \emph{2022 ACM Conference on Fairness, Accountability, and
  Transparency}, pages 2047--2056.

\bibitem[{Bahl et~al.(1983)Bahl, Jelinek, and Mercer}]{bahl1983maximum}
Lalit~R Bahl, Frederick Jelinek, and Robert~L Mercer. 1983.
\newblock A maximum likelihood approach to continuous speech recognition.
\newblock \emph{IEEE transactions on pattern analysis and machine
  intelligence}, (2):179--190.

\bibitem[{Baker(1975{\natexlab{a}})}]{baker1975dragon}
James Baker. 1975{\natexlab{a}}.
\newblock The dragon system--an overview.
\newblock \emph{IEEE Transactions on Acoustics, speech, and signal Processing},
  23(1):24--29.

\bibitem[{Baker(1975{\natexlab{b}})}]{baker1975auto}
James~K. Baker. 1975{\natexlab{b}}.
\newblock \emph{Stochastic Modeling as a Means of Automatic Speech
  Recognition.}
\newblock Ph.D. thesis, USA.

\bibitem[{Buckner(2018)}]{buckner2018empiricism}
Cameron Buckner. 2018.
\newblock Empiricism without magic: Transformational abstraction in deep
  convolutional neural networks.
\newblock \emph{Synthese}, 195(12):5339--5372.

\bibitem[{Buckner(2019)}]{buckner2019deep}
Cameron Buckner. 2019.
\newblock Deep learning: A philosophical introduction.
\newblock \emph{Philosophy compass}, 14(10):e12625.

\bibitem[{Cahyawijaya et~al.(2022)Cahyawijaya, Aji, Lovenia, Winata, Wilie,
  Mahendra, Koto, Moeljadi, Vincentio, Romadhony
  et~al.}]{cahyawijaya2022nusacrowd}
Samuel Cahyawijaya, Alham~Fikri Aji, Holy Lovenia, Genta~Indra Winata, Bryan
  Wilie, Rahmad Mahendra, Fajri Koto, David Moeljadi, Karissa Vincentio, Ade
  Romadhony, et~al. 2022.
\newblock Nusacrowd: A call for open and reproducible nlp research in
  indonesian languages.
\newblock \emph{arXiv preprint arXiv:2207.10524}.

\bibitem[{Celikyilmaz et~al.(2018)Celikyilmaz, Deng, and
  Hakkani-T{\"u}r}]{celikyilmaz2018deep}
Asli Celikyilmaz, Li~Deng, and Dilek Hakkani-T{\"u}r. 2018.
\newblock Deep learning in spoken and text-based dialog systems.
\newblock In \emph{Deep Learning in Natural Language Processing}, pages 49--78.
  Springer.

\bibitem[{Chang et~al.(2020)Chang, Rong, Xu, Huang, Zhang, Cui, Zhu, and
  Huang}]{chang2020restricted}
Heng Chang, Yu~Rong, Tingyang Xu, Wenbing Huang, Honglei Zhang, Peng Cui, Wenwu
  Zhu, and Junzhou Huang. 2020.
\newblock A restricted black-box adversarial framework towards attacking graph
  embedding models.
\newblock In \emph{Proceedings of the AAAI Conference on Artificial
  Intelligence}, volume~34, pages 3389--3396.

\bibitem[{Chen et~al.(2021)Chen, Tworek, Jun, Yuan, Pinto, Kaplan, Edwards,
  Burda, Joseph, Brockman et~al.}]{chen2021evaluating}
Mark Chen, Jerry Tworek, Heewoo Jun, Qiming Yuan, Henrique Ponde de~Oliveira
  Pinto, Jared Kaplan, Harri Edwards, Yuri Burda, Nicholas Joseph, Greg
  Brockman, et~al. 2021.
\newblock Evaluating large language models trained on code.
\newblock \emph{arXiv preprint arXiv:2107.03374}.

\bibitem[{Chiticariu et~al.(2013)Chiticariu, Li, and
  Reiss}]{chiticariu2013rule}
Laura Chiticariu, Yunyao Li, and Frederick Reiss. 2013.
\newblock Rule-based information extraction is dead! long live rule-based
  information extraction systems!
\newblock In \emph{Proceedings of the 2013 conference on empirical methods in
  natural language processing}, pages 827--832.

\bibitem[{Chomsky(1956{\natexlab{a}})}]{Chomsky1956logical}
Noam Chomsky. 1956{\natexlab{a}}.
\newblock The logical structure of linguistic theory.
\newblock \emph{Synthese}, 40(2):317--352.

\bibitem[{Chomsky(1956{\natexlab{b}})}]{chomsky1956three}
Noam Chomsky. 1956{\natexlab{b}}.
\newblock Three models for the description of language.
\newblock \emph{IRE Transactions on information theory}, 2(3):113--124.

\bibitem[{Chomsky(1957)}]{chomsky1957syntactic}
Noam Chomsky. 1957.
\newblock Syntactic structures.
\newblock In \emph{Syntactic Structures}. De Gruyter Mouton.

\bibitem[{Chomsky(1963)}]{chomsky1963formal}
Noam Chomsky. 1963.
\newblock Formal properties of grammars.
\newblock \emph{Handbook of Math. Psychology}, 2:328--418.

\bibitem[{Chomsky(1965)}]{chomsky1965aspects}
Noam Chomsky. 1965.
\newblock Aspects of the theory of syntax.
\newblock \emph{Multilingual Matters: MIT Press}.

\bibitem[{Chomsky and Halle(1968)}]{chomsky1968sound}
Noam Chomsky and Morris Halle. 1968.
\newblock The sound pattern of english.

\bibitem[{Collobert et~al.(2011)Collobert, Weston, Bottou, Karlen, Kavukcuoglu,
  and Kuksa}]{collobert2011natural}
Ronan Collobert, Jason Weston, L{\'e}on Bottou, Michael Karlen, Koray
  Kavukcuoglu, and Pavel Kuksa. 2011.
\newblock Natural language processing (almost) from scratch.
\newblock \emph{Journal of machine learning research}, 12(ARTICLE):2493--2537.

\bibitem[{Crane(2018)}]{crane2018questionable}
Matt Crane. 2018.
\newblock Questionable answers in question answering research: Reproducibility
  and variability of published results.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  6:241--252.

\bibitem[{de~Freitas~Pereira et~al.(2022)de~Freitas~Pereira, Schmidli, Linghu,
  Zhang, Marcel, and G{\"u}nther}]{de2022eight}
Tiago de~Freitas~Pereira, Dominic Schmidli, Yu~Linghu, Xinyi Zhang,
  S{\'e}bastien Marcel, and Manuel G{\"u}nther. 2022.
\newblock Eight years of face recognition research: Reproducibility,
  achievements and open issues.
\newblock \emph{arXiv}, (2208.04040).

\bibitem[{Fadel et~al.(2019)Fadel, Tuffaha, Al-Ayyoub et~al.}]{fadel2019arabic}
Ali Fadel, Ibraheem Tuffaha, Mahmoud Al-Ayyoub, et~al. 2019.
\newblock Arabic text diacritization using deep neural networks.
\newblock In \emph{2019 2nd international conference on computer applications
  \& information security (ICCAIS)}, pages 1--7. IEEE.

\bibitem[{Farooq and Hafeez(2020)}]{farooq2020covid}
Muhammad Farooq and Abdul Hafeez. 2020.
\newblock Covid-resnet: A deep learning framework for screening of covid19 from
  radiographs.
\newblock \emph{arXiv preprint arXiv:2003.14395}.

\bibitem[{Frank et~al.(2019)Frank, Monaghan, and Tsoukala}]{frank2019neural}
Stefan~L Frank, Padraic Monaghan, and Chara Tsoukala. 2019.
\newblock Neural network models of language acquisition and processing.
\newblock In \emph{Human language: From genes and brain to behavior}, pages
  277--293. MIT Press.

\bibitem[{Gibson and Cano(2022)}]{gibson2022productive}
Perry Gibson and Jos{\'e} Cano. 2022.
\newblock Productive reproducible workflows for dnns: A case study for
  industrial defect detection.
\newblock \emph{arXiv preprint arXiv:2206.09359}.

\bibitem[{Guo et~al.(2022)Guo, Li, Jiang, and Liu}]{guo2022freetransfer}
Yinpeng Guo, Liangyou Li, Xin Jiang, and Qun Liu. 2022.
\newblock Freetransfer-x: Safe and label-free cross-lingual transfer from
  off-the-shelf models.
\newblock \emph{arXiv preprint arXiv:2206.06586}.

\bibitem[{He and Deng(2018)}]{he2018deep}
Xiaodong He and Li~Deng. 2018.
\newblock Deep learning in natural language generation from images.
\newblock In \emph{Deep Learning in Natural Language Processing}, pages
  289--307. Springer.

\bibitem[{Humboldt(1836)}]{humboldt1836uber}
W~von Humboldt. 1836.
\newblock Uber die verschiedenheit des menschlichen sprachbaues.
\newblock \emph{Berlin: Konigliche Akademie der Wissenschaften}.

\bibitem[{Jelinek et~al.(1990)Jelinek, Merialdo, Roukos, Strauss
  et~al.}]{jelinek1990self}
Fred Jelinek, B~Merialdo, S~Roukos, M~Strauss, et~al. 1990.
\newblock Self-organized language modeling for speech recognition.
\newblock In \emph{Readings in speech recognition}.

\bibitem[{Jelinek(1976)}]{jelinek1976continuous}
Frederick Jelinek. 1976.
\newblock Continuous speech recognition by statistical methods.
\newblock \emph{Proceedings of the IEEE}, 64(4):532--556.

\bibitem[{Jurafsky and Martin(2022)}]{jurafsky2022speech}
Daniel Jurafsky and James~H Martin. 2022.
\newblock Speech and language processing: An introduction to natural language
  processing, computational linguistics, and speech recognition (draft of
  january 12, 2022).

\bibitem[{Kapoor and Narayanan(2022)}]{kapoor2022leakage}
Sayash Kapoor and Arvind Narayanan. 2022.
\newblock Leakage and the reproducibility crisis in ml-based science.
\newblock \emph{arXiv preprint arXiv:2207.07048}.

\bibitem[{Klein et~al.(2017)Klein, Kim, Deng, Senellart, and
  Rush}]{klein2017opennmt}
Guillaume Klein, Yoon Kim, Yuntian Deng, Jean Senellart, and Alexander~M Rush.
  2017.
\newblock Opennmt: Open-source toolkit for neural machine translation.
\newblock \emph{arXiv preprint arXiv:1701.02810}.

\bibitem[{Laurence and Margolis(2001)}]{laurence2001poverty}
Stephen Laurence and Eric Margolis. 2001.
\newblock The poverty of the stimulus argument.
\newblock \emph{British Journal for the Philosophy of Science}, 52(2).

\bibitem[{LeCun et~al.(2015)LeCun, Bengio, Hinton et~al.}]{lecun2015deep}
Yann LeCun, Yoshua Bengio, Geoffrey Hinton, et~al. 2015.
\newblock Deep learning.
\newblock \emph{Nature, 521 (7553), 436-444}.

\bibitem[{Manning and Schutze(1999)}]{manning1999foundations}
Christopher Manning and Hinrich Schutze. 1999.
\newblock \emph{Foundations of statistical natural language processing}.
\newblock MIT press.

\bibitem[{Marcus(2018)}]{marcus2018deep}
Gary Marcus. 2018.
\newblock Deep learning: A critical appraisal.
\newblock \emph{arXiv preprint arXiv:1801.00631}.

\bibitem[{Markov(2006)}]{markov2006example}
Andre{\u\i}~Andreevich Markov. 2006.
\newblock An example of statistical investigation of the text eugene onegin
  concerning the connection of samples in chains.
\newblock \emph{Science in Context}, 19(4):591--600.

\bibitem[{Miller and Chomsky(1963)}]{miller1963finitary}
George~A Miller and Noam Chomsky. 1963.
\newblock Finitary models of language users. handbook of mathematical
  psychology, vol. by r. duncan luce, robert r. bush, and eugene galanter,
  419--91.

\bibitem[{Mnih and Hinton(2008)}]{mnih2008scalable}
Andriy Mnih and Geoffrey~E Hinton. 2008.
\newblock A scalable hierarchical distributed language model.
\newblock \emph{Advances in neural information processing systems}, 21.

\bibitem[{Morin and Bengio(2005)}]{morin2005hierarchical}
Frederic Morin and Yoshua Bengio. 2005.
\newblock Hierarchical probabilistic neural network language model.
\newblock In \emph{International workshop on artificial intelligence and
  statistics}, pages 246--252. PMLR.

\bibitem[{Russell and Norvig(2021)}]{russell2021artificial}
Stuart Russell and Peter Norvig. 2021.
\newblock \emph{Artificial Intelligence: A Modern Approach}.
\newblock Pearson.

\bibitem[{Schramowski et~al.(2022)Schramowski, Turan, Andersen, Rothkopf, and
  Kersting}]{schramowski2022large}
Patrick Schramowski, Cigdem Turan, Nico Andersen, Constantin~A Rothkopf, and
  Kristian Kersting. 2022.
\newblock Large pre-trained language models contain human-like biases of what
  is right and wrong to do.
\newblock \emph{Nature Machine Intelligence}, 4(3):258--268.

\bibitem[{Shannon(1948)}]{shannon1948mathematical}
Claude~Elwood Shannon. 1948.
\newblock A mathematical theory of communication.
\newblock \emph{The Bell system technical journal}, 27(3):379--423.

\bibitem[{Siegle et~al.(2015)Siegle, Hale, Newman, and
  Voigts}]{siegle2015neural}
Joshua~H Siegle, Gregory~J Hale, Jonathan~P Newman, and Jakob Voigts. 2015.
\newblock Neural ensemble communities: open-source approaches to hardware for
  large-scale electrophysiology.
\newblock \emph{Current opinion in neurobiology}, 32:53--59.

\bibitem[{Silva et~al.(2022)Silva, Siqueira, Tarrega, Beinotti, Nunes, Gardini,
  da~Silva, da~Silva, and de~Carvalho}]{silva2022no}
Mar{\'\i}lia Costa~Rosendo Silva, Felipe~Alves Siqueira, Jo{\~a}o
  Pedro~Mantovani Tarrega, Jo{\~a}o Vitor~Pataca Beinotti, Augusto~Sousa Nunes,
  Miguel de~Mattos Gardini, Vin{\'\i}cius Adolfo~Pereira da~Silva, N{\'a}dia
  F{\'e}lix~Felipe da~Silva, and Andr{\'e} Carlos Ponce de Leon~Ferreira
  de~Carvalho. 2022.
\newblock No pattern, no recognition: a survey about reproducibility and
  distortion issues of text clustering and topic modeling.
\newblock \emph{arXiv preprint arXiv:2208.01712}.

\bibitem[{Silver et~al.(2017)Silver, Schrittwieser, Simonyan, Antonoglou,
  Huang, Guez, Hubert, Baker, Lai, Bolton et~al.}]{silver2017mastering}
David Silver, Julian Schrittwieser, Karen Simonyan, Ioannis Antonoglou, Aja
  Huang, Arthur Guez, Thomas Hubert, Lucas Baker, Matthew Lai, Adrian Bolton,
  et~al. 2017.
\newblock Mastering the game of go without human knowledge.
\newblock \emph{nature}, 550(7676):354--359.

\bibitem[{Srivastava et~al.(2022)Srivastava, Rastogi, Rao, Shoeb, Abid, Fisch,
  Brown, Santoro, Gupta, Garriga-Alonso et~al.}]{srivastava2022beyond}
Aarohi Srivastava, Abhinav Rastogi, Abhishek Rao, Abu Awal~Md Shoeb, Abubakar
  Abid, Adam Fisch, Adam~R Brown, Adam Santoro, Aditya Gupta, Adri{\`a}
  Garriga-Alonso, et~al. 2022.
\newblock Beyond the imitation game: Quantifying and extrapolating the
  capabilities of language models.
\newblock \emph{arXiv preprint arXiv:2206.04615}.

\bibitem[{Stich(1978)}]{stich1978empiricism}
Stephen~P Stich. 1978.
\newblock Empiricism, innateness, and linguistic universals.
\newblock \emph{Philosophical Studies: An International Journal for Philosophy
  in the Analytic Tradition}, 33(3):273--286.

\bibitem[{Suresha et~al.(2018)Suresha, Kidzi{\'n}ski, Halilaj, Gold, and
  Delp}]{suresha2018automated}
S~Suresha, L~Kidzi{\'n}ski, E~Halilaj, GE~Gold, and SL~Delp. 2018.
\newblock Automated staging of knee osteoarthritis severity using deep neural
  networks.
\newblock \emph{Osteoarthritis and Cartilage}, 26:S441.

\bibitem[{Towers et~al.(2022)Towers, Forshaw, Atapour-Abarghouei, and
  McGough}]{towers2022long}
David Towers, Matthew Forshaw, Amir Atapour-Abarghouei, and Andrew~Stephen
  McGough. 2022.
\newblock Long-term reproducibility for neural architecture search.
\newblock \emph{arXiv preprint arXiv:2207.04821}.

\bibitem[{Tur et~al.(2018)Tur, Celikyilmaz, He, Hakkani-T{\"u}r, and
  Deng}]{tur2018deep}
Gokhan Tur, Asli Celikyilmaz, Xiaodong He, Dilek Hakkani-T{\"u}r, and Li~Deng.
  2018.
\newblock Deep learning in conversational language understanding.
\newblock In \emph{Deep Learning in Natural Language Processing}, pages 23--48.
  Springer.

\bibitem[{Weidinger et~al.(2022)Weidinger, Uesato, Rauh, Griffin, Huang,
  Mellor, Glaese, Cheng, Balle, Kasirzadeh et~al.}]{weidinger2022taxonomy}
Laura Weidinger, Jonathan Uesato, Maribeth Rauh, Conor Griffin, Po-Sen Huang,
  John Mellor, Amelia Glaese, Myra Cheng, Borja Balle, Atoosa Kasirzadeh,
  et~al. 2022.
\newblock Taxonomy of risks posed by language models.
\newblock In \emph{2022 ACM Conference on Fairness, Accountability, and
  Transparency}, pages 214--229.

\bibitem[{Wieling et~al.(2018)Wieling, Rawee, and van
  Noord}]{wieling2018reproducibility}
Martijn Wieling, Josine Rawee, and Gertjan van Noord. 2018.
\newblock Reproducibility in computational linguistics: are we willing to
  share?
\newblock \emph{Computational Linguistics}, 44(4):641--649.

\bibitem[{Xu et~al.(2022)Xu, Alon, Neubig, and Hellendoorn}]{xu2022systematic}
Frank~F Xu, Uri Alon, Graham Neubig, and Vincent~Josua Hellendoorn. 2022.
\newblock A systematic evaluation of large language models of code.
\newblock In \emph{Proceedings of the 6th ACM SIGPLAN International Symposium
  on Machine Programming}, pages 1--10.

\bibitem[{y~Arcas(2022)}]{y2022large}
Blaise~Ag{\"u}era y~Arcas. 2022.
\newblock Do large language models understand us?
\newblock \emph{Daedalus}, 151(2):183--197.

\bibitem[{Zuin et~al.(2020)Zuin, Veloso, Portinari, and
  Ziviani}]{zuin2020automatic}
Gianluca Zuin, Adriano Veloso, Jo{\~a}o~C{\^a}ndido Portinari, and Nivio
  Ziviani. 2020.
\newblock Automatic tag recommendation for painting artworks using diachronic
  descriptions.
\newblock In \emph{2020 International Joint Conference on Neural Networks
  (IJCNN)}, pages 1--8. IEEE.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt]{article}

\usepackage[]{emnlp2021}

\usepackage{times}
\usepackage{latexsym}

\usepackage[T1]{fontenc}

\usepackage[utf8]{inputenc}

\usepackage{microtype}

\title{Inaccessible Neural Language Models Could \\ Reinvigorate Linguistic Nativism}

\author{Patrick Perrine  \\
  California Polytechnic State University \\
  1 Grand Ave, San Luis Obispo, CA, 93410 \\
  \texttt{paperrin@calpoly.edu}}
  
\begin{document}
\maketitle


\section*{Abstract}
Large Language Models (LLMs) have been making big waves in the machine learning community within the past few years. The impressive scalability of LLMs due to the advent of deep learning can be seen as a continuation of empiricist lingusitic methods, as opposed to rule-based linguistic methods that are grounded in a nativist perspective. Current LLMs are generally inaccessible to resource-constrained researchers, due to a variety of factors including closed source code. This work argues that this lack of accessibility could instill a nativist bias in researchers new to computational linguistics, given that new researchers may only have rule-based, nativist approaches to study to produce new work. Also, given that there are numerous critics of deep learning claiming that LLMs and related methods may soon lose their relevancy, we speculate that such an event could trigger a new wave of nativism in the language processing community. To prevent such a dramatic shift and placing favor in hybrid methods of rules and deep learning, we call upon researchers to open source their LLM code wherever possible to allow both empircist and hybrid approaches to remain accessible.

\section{Introduction}

Large Language Models (LLMs) have been a popular topic of research among the academic community \cite{srivastava2022beyond}. The promise of a near-general purpose neural model for a variety of language processing tasks is indeed an attractive one \cite{xu2022systematic}. Deep learning has made significant developments in language tasks such as conversational language understanding \cite{tur2018deep}, spoken/text-based dialog systems \cite{celikyilmaz2018deep}, and natural language generation from images \cite{he2018deep}. Large Language models can be viewed as the natural progression away from the rigid rule-based systems that we've had since the 1950's \cite{chiticariu2013rule}, continuing the empiricist mentality of statistical natural language processing without the potentially costly and context-specific activity of feature engineering \cite{collobert2011natural}. However, with large corporations touting their ever-growing, state-of-the-art models under closed-source code and payment walls, it could be seen that these large language models are becoming \emph{less accessible}. Some organizations have acknowledged the potential harms that deep learning models could cause by establishing ethical frameworks \cite{ashurst2022ai} \cite{weidinger2022taxonomy}, but there are still growing concerns regarding accessibility and the result of false/irreproducible science \cite{kapoor2022leakage}. 

This criticism for empiricist methods is not new in linguistics-based science, in that Chomsky's Poverty of the Stimulus Argument \cite{stich1978empiricism} has a rich history of discussion and debate amongst linguists, scientists, and philosophers \cite{laurence2001poverty}. In this work, we will briefly introduce this debate over language learning between nativists and empiricists, relate these topics to research in natural language processing, and disucss how the current state of this research is reinforcing an imbalance between the two perspectives. We intend to deliver a neutral ground of analysis, as we agree that a hybrid approach for NLP research can lead to strong results. The current bias towards the highly-popular, but inaccessible empiricist methods utilizing LLMs could lead to a new wave of nativism in natural language processing work, following a large backlash against such empirical methods.

\section{Background}
We now provide a holistic background on the linguistic and scientific developments that encompass this issue.

\subsection{The Three Waves of Modern NLP}

We will give a brief background on the three main waves of modern natural language processing research: the rule-based theories popularized by Noam Chomsky \cite{chomsky1965aspects}, the statistics-based empiricist experiments \cite{jelinek1976continuous}, and today's popular methodology of deep learning for natural language processing \cite{collobert2011natural}. The first wave is considered to be under a nativist perspective \cite{laurence2001poverty}, whereas the latter waves are in support of an empiricist lens \cite{frank2019neural}.

\subsubsection{Rule-based NLP}

The concept of viewing language as a static system of rules to determine interpretation has been present as early as the 1830's \cite{humboldt1836uber}. Noam Chomsky popularized this perspective in the domain of linguistics as a challenge to an existing overbearance of empiricist methods \cite{chomsky1956three, laurence2001poverty}. 

This rule-based approach to linguistics dominated the field for decades, following Chomsky's mutliple works emphasizing and reinforcing this doctrine \cite{Chomsky1956logical, chomsky1957syntactic, chomsky1963formal, chomsky1965aspects, chomsky1968sound}. Being based in propositional logic and a fixed content, rule-based methods are arguably rather accessible to researchers with limited resources. These methods continued to be prevalent in the field until the 1970's, when statistical methods were proven to be very useful.

\subsubsection{Statistical NLP}

The roots of statistical language processing stem from Andrey Markov's efforts in computing bigram and trigram probabilities \cite{jurafsky2022speech} of vowel/consonant predictions using a novel as a corpus in 1913 \cite{markov2006example}. This $n$-gram approach was later applied to predicting sequences of English words \cite{shannon1948mathematical}. This popularized the notion of using Markov chains for use in a variety of applications within and outside of linguistics. 

Chomsky specifically challenged this use of finite-state Markov processes, the processes that formed $n$-gram based approaches, to be useless in serving as a comprehensive cognitive model of grammatical knowledge in humans \cite{chomsky1956three, chomsky1957syntactic, miller1963finitary}. This hindered the progress of probabilistic approaches in linguistics.

Over a decade later, statistical language processing was revitalized due in part to a series of successful experiments using $n$-gram models for speech recognition \cite{baker1975dragon, baker1975auto, jelinek1976continuous, bahl1983maximum, jelinek1990self}. These empiricist-based experiments showed that Chomsky's nativist theories do not extend to recognizing speech in real time as previously proposed \cite{chomsky1968sound}.

This marked a shift towards looking at language processing through an empirical lens, where a hypothesis test primarily guides the experimentation process, rather than theoretical insights \cite{manning1999foundations}. After the successful statistical speech recognition experiments of the mid 1970's, statistical NLP reigned as the dominant approach for decades. 

\subsubsection{ML-based NLP}
Researchers soon began to use shallow neural networks to reinforce statistical methodologies in NLP. In the late 2000's, the advent of deeper neural networks for NLP began to stir when scalable, hierarchical language models \cite{morin2005hierarchical, mnih2008scalable} and increased computing power became available for use by researchers. 

Alongside these developments, researchers became tiresome of having to hand-engineer features for neural networks to learn from, as this can be a costly and rather context-specific task \cite{collobert2011natural}. In was in the 2010's that deep learning became known more globally \cite{lecun2015deep}, with NLP being a highly prominent application for deep neural networks. This sparked the current practice of training large language models in efforts to create a general model for many language tasks \cite{srivastava2022beyond}. In essence, the empiricist era of NLP has persisted to today through the evolution of deep learning practices. Some applications of deep learning outside of language have even used empiricist terms such as \emph{tabula rasa} very openly \cite{silver2017mastering}. The use of deep neural networks for language tasks has been confirmed to reinforce empircist ideology \cite{frank2019neural}.

\section{Deep Learning Can Be Inaccessible}

Deep learning as a science has been under fire for a number of reasons. While there have been encouraging results across many application domains of deep learning and positive insights about their role in advancing empiricism \cite{buckner2018empiricism}, deep learning has garnered skepticsm from both in and outside of its community \cite{marcus2018deep, buckner2019deep}. 

These critcisms of deep NLP can stem from a \emph{lack of open sourcing of model code} and also data \cite{klein2017opennmt, fadel2019arabic, chen2021evaluating, guo2022freetransfer, xu2022systematic}. These issues are not exclusive to language processing, as other domains have reasons to leave aspects of their experimentation private or inaccessible when publishing \cite{siegle2015neural, suresha2018automated, farooq2020covid, zuin2020automatic, guo2022freetransfer}.

We now focus on issues with closed-source large language models due to their popularity and the recent claims of greater intelligence (even sentience), as opposed to other models \cite{y2022large}.

\section{Potential Harms}

\subsection{Potential Harms of Open-Sourcing LLMs}

To offer a well-rounded argument in favor of open-sourcing LLMs, we will briefly cover some intuitions behind close-sourcing them in terms of potential harms. 

LLMs could be repurposed for malicious purposes, particularly in generative tasks. LLMs have been seen to learn negative human biases/patterns in speech such as hate speech, discrimination, and the promotion of misinformation \cite{schramowski2022large}. If a powerful, pre-trained LLM is made open source, then it could be repurposed as an engine to cause harm across the internet at great scale \cite{weidinger2022taxonomy}. It could also be argued that open sourcing LLM code that has been deployed to end-users could pose security risks \cite{chang2020restricted}.

We counter the argument of potential LLM misuse by malicious parties by arguing that such models or derivatives of such should not be published in any form, open or closed source. We argue that LLM experimental papers that indicate such potential to cause harm at scale should be filtered out at the publication review stage, something that has been discussed in the deep learning community as of late \cite{ashurst2022ai}. We also counter the security concern argument by saying that this could hold true for all open source software that is deployable, not just LLMs.

\subsection{Potential Harms from Continued Close-Sourcing of LLMs}

We argue that there are more potential harms in the continued prevalence of close sourced LLM code than the potential harms of open sourcing them.

\subsubsection{Nativist Biases}

Given that LLM experiments are becoming so large, costly, and complex, it is difficult to argue that an independent researcher can stake a claim in this subfield. With top publication venues focusing heavily on empiricist experimentation \cite{russell2021artificial}, researchers outside the typical corporate scope of research could be incentivized to explore nativist, rule-based approaches to solve problems in the NLP domain. If it is the empiricist group's better interest to foster growth in their methodologies and not opposing methods, steps should be taken in order to make their approaches accessible. Also, for hybrid methods to function, an ML-based solution should be made accessible to combine with the ruleset from the nativist side. This trend could be fostering a new generation of Chomsky-following nativist NLP researchers, which would not bode well for empiricists if the public begins to lose interest in deep learning methods for NLP.

\subsubsection{Lack of Reproducibility}

We mention reproducibility and will further clarify its meaning due to an also recent, yet broader problem in deep learning research, the reproducibility crisis \cite{kapoor2022leakage}. Not only are large language models becoming difficult to reproduce, results from other areas of ML are becoming difficult to produce \cite{de2022eight, towers2022long}. Initiatives to measure reproducibility across publication venues have been created, such as the \href{https://www.cs.mcgill.ca/~jpineau/ICLR2018-ReproducibilityChallenge.html}{ML Reproducibility Challenge}. LLM experiments have been specifically reviewed to have a questionable about of reproducibility \cite{crane2018questionable, wieling2018reproducibility, cahyawijaya2022nusacrowd, silva2022no}. There is also implied to be a significant amount of computational irreproducibility of LLM experimentation, given model complexity and data, however, we leave this exploration for future work.

There is some hope in the form of positive reproducibility reports in deep learning \cite{gibson2022productive}. However, this growing amount of ``bad press'' for deep learning, specifically LLMs, could cause the public to begin distrusting LLM research. This, again, could trigger a revisiting of Chomsky's rule-based theories of language.

\subsubsection{Issues in NLP Education}

Given the previously mentioned issues, this lack of accessibility could affect the education of NLP methods. If students do not have access to code of LLMs, it could be difficult for them to learn to implement complex language model code of their own and learn to keep up with the state of the art. A lack of reproducibility could also be disenfranchising to a young, empircist NLP researcher, leading them to pursue nativist approaches. These issues could reinforce the use of statistical, pre-deep learning techniques in the classroom, but it is difficult to argue that publication venues are interested in shallow neural network experimentation at this time.

These issues combine to form an uneven playing field for students to study NLP in empiricist and hybrid forms. After studying NLP formally, they may be inclined to commit to nativist methods or even reinforce the popularity of them at scale.

\section{Potential Solution}

We ask that publication venues merit open source LLM experiments significantly higher than they do currently. We believe that this would mitigate the issues discussed previously in this work. There seem to be developments occuring now in the deep learning publication space to help implement this in a proper form of governance \cite{ashurst2022ai}.

\section{Conclusion}
In this work, we provided a comprehensive history of natural language processing methodologies over roughly the past century. We then used this narrative to lead into today's deep learning practices used in language processing, and current issues in an excessive closed sourcing of code for LLMs. It is our hope that this work inspires researchers and reviewers to champion open source language model code in order to pave the way for a more balanced research space.

\bibliographystyle{acl_natbib}
\bibliography{custom}

\end{document}